# All-in-One Slider for Attribute Manipulation in Diffusion Models

### v1

@@ 2    Additional Experimental Results
We present additional qualitative results demonstrating our method’s ability to manipulate a wide range of facial attributes in an identity-preserving manner, as illustrated in Figures 11 to 20. Each figure below shows a series of faces in rows, where the leftmost column is the original input image, and the rightward images represent progressive edits along a specific semantic attribute.
3    Prompt Information
@@
For reproducibility, we provide the text prompts and associated random seeds used to generate the images in the appendix figures.
Prompts for Main Figures
@@
Below are the prompts used to generate the images shown in the main figures of the paper.
@@
Figure 1: (1) A photo of a gentle Thai man with layered hair, in a warm-toned studio. 33094
@@
Figure 1: (2) A photo of a honest Dutch man in a white button-up shirt, with a blurred background. 27454
@@
Figure 1: (3) A high-quality portrait of a man in a tuxedo, with clear, detailed facial features. 33673
@@
Figure 4: Prompt: A close-up of a woman wearing a red scarf, looking thoughtful. 31796
Figure 5: A photo of a honest Latin American talented woman. 33819
@@
Figure 6: (1) A woman with medium-length blonde hair, pale skin, and soft blue eyes.33819
Figure 6: (2) A photo of a peaceful Middle Eastern woman. 22436
@@
Figure 8: A photo of a woman wearing a red scarf. 42
Prompts for Appendix Figures
Below are the prompts used to generate the images shown in the appendix.
@@
(1) A photo of a charming Singaporean man with clean-cut hair, with studio backdrop. 16727
(2) A photo of a gentle Thai man with layered hair, in a warm-toned studio. 26662
(3) A photo of a composed Swedish man with short blond hair, with a blurred background. 38836
(4) A photo of a gentle Vietnamese man with short cropped hair, with a soft gray backdrop. 20664
(5) A photo of a peaceful Middle Eastern woman. 42
(6) A photo of a composed Moroccan man wearing a woolen scarf, with a blurred background. 29425
(7) A photo of a kind Turkish adult woman. 26281
(8) A photo of a honest Dutch man in a white button-up shirt, with a blurred background. 32119
@@
(1) A man in a white button-up shirt, with soft natural lighting on his face. 3203
(2) A photo of a refined Italian man wearing a checkered blazer, with soft window light. 22088
(3) A photo of a calm South African man wearing round glasses, with a soft gray backdrop. 6839
(4) A photo of a radiant American man wearing a casual sweater, with elegant lighting. 28031
(5) A photo of a gentle Vietnamese man with short cropped hair, with a soft gray backdrop. 46112
(6) A photo of a serene Danish man with round glasses, near a sunlit window. 37614
(7) A photo of a poised Greek person in a linen shirt, in a studio with soft shadows. 7242
(8) A photo of a thoughtful Latin American woman in a white button-up shirt, with a blurred background. 26731
@@
(1) A photo of a stylish Serbian person with long hair, in a candle-lit room. 42
(2) A photo of a refined Italian man wearing a checkered blazer, with soft window light. 48991
(3) A photo of a graceful British man with neatly combed hair, in a bright workspace. 33898
(4) A photo of a composed Swedish man with short blond hair, with a blurred background. 38623
(5) A photo of a stylish Belgian person with a fringe haircut, in a softly lit room. 44007
(6) A photo of a poised Greek person in a linen shirt, in a studio with soft shadows. 10133
(7) A photo of a poised Greek person in a linen shirt, in a studio with soft shadows. 7242
(8) A photo of a composed Swedish man with short blond hair, with a blurred background. 38836
Figure 14 prompts:
(1) A photo of a honest Latin American talented woman. 33819
(2) A photo of a contemplative Pakistani woman with a patterned scarf, with natural daylight. 46948
(3) A photo of a charming Mongolian person wearing a fur-lined coat, in snowy weather. 2337
(4) A photo of a sophisticated Russian woman with a pearl necklace, in a minimalist room. 6603
(5) A photo of a contemplative Pakistani woman with a patterned scarf, with natural daylight. 26926
(6) A photo of a unique Filipino woman. 48084
(7) A woman with short hair, looking at the camera. 13476
(8) A photo of a charming Taiwanese person with medium-length hair, in a sunlit studio. 26595
Figure 15 prompts:
@@
(1) A portrait of a woman standing by the beach at sunset, with soft and natural lighting. 42
(2) A photo of a woman smiling in a cafe setting. 42
(3) A photo of a elegant and kind woman. 344980
(4) A photo of a beautiful woman standing by the beach at sunset. 42
(5) A woman with short hair, looking at the camera. 39743
(6) A photo of a radiant American man wearing a casual sweater, with elegant lighting. 28031
(7) A photo of a honest Polish woman in a white button-up shirt, with a soft gray backdrop. 29600
(8) A photo of a thoughtful Latin American woman in a white button-up shirt, with a blurred background. 26731
Figure 16 prompts:
(1) A photo of a serene Danish man with round glasses, near a sunlit window. 37614
(2) A photo of a serene Danish man with round glasses, near a sunlit window. 47741
(3) A photo of a gentle Vietnamese man with short cropped hair, with a soft gray backdrop. 26529
(4) A photo of a gentle Vietnamese man with short cropped hair, with a soft gray backdrop. 46112
(5) A photo of a stylish Serbian person with long hair, in a candle-lit room. 42
(6) A photo of a composed Swedish man with short blond hair, with a blurred background. 29070
(7) A photo of a gentle Vietnamese man with short cropped hair, with a soft gray backdrop. 19257
(8) A photo of a gentle Vietnamese man with short cropped hair, with a soft gray backdrop. 26125
Figure 17 prompts:
(1) A portrait of a woman standing by the beach at sunset, with soft and natural lighting. 42
(2) A woman with short hair, looking at the camera. 13476
(3) A woman with short hair, looking at the camera. 41815
(4) A woman with short hair, looking at the camera. 39743
(5) A photo of a sophisticated Russian woman with a pearl necklace, in a minimalist room. 21553
(6) A photo of a sophisticated Russian woman with a pearl necklace, in a minimalist room. 6603
Figure 18: prompts:
(1) A photo of a honest Dutch man in a white button-up shirt, with elegant lighting. 43448
(2) A photo of a elegant Chinese man wearing a red scarf, with a blurred background. 26099
(3) A photo of a confident Turkish man wearing a floral crown, with natural daylight. 2606
(4) A photo of a cheerful Filipino man wearing a red scarf, with a soft gray backdrop. 21968
(5) A woman with short hair, looking at the camera. 13476
(6) A photo of a vibrant Indian man. 41026
(7) A photo of a serene American man wearing a floral crown, with a soft gray backdrop. 41682
(8) A photo of a honest Middle Eastern adult woman wearing a red scarf, in a studio. 49637
Figure 19 prompts:
(1) A photo of a beautiful woman. 45678
(2) A photo of a beautiful woman. 355667
(3) A photo of a wise European man wearing a casual sweater, in a sunny environment. 26509
(4) A photo of a elegant woman. 569988
Figure 20 prompts:
(1) A portrait of a woman standing by the beach at sunset, with soft and natural lighting. 42
(2) A close-up of a woman with a floral crown on her head. 43533
(3) A photo of a honest Latin American talented woman. 33819
(4) A close-up portrait of a indian woman. 1646
(5) A photo of a radiant Nepalese woman wearing a gold necklace, in natural light. 26268
(6) A photo of a gentle Thai man with layered hair, in a warm-toned studio. 31527
(7) A photo of a gentle Vietnamese man with short cropped hair, with a soft gray backdrop. 10368
(8) A photo of a serene Brazilian confident woman. 45433
Note that the prompts correspond to the images in order, starting from the top-left image and proceeding row by row (left to right, then top to bottom).
@@
Figure 11: Continuous manipulation of the Smile attribute. Each row shows a subject transitioning from neutral to smiling, with identity and lighting well preserved.
Figure 12: Continuous manipulation of the Old attribute. Each subject’s image is progressively transformed to appear older, showing realistic aging effects such as wrinkles and hair graying, while maintaining overall identity.
Figure 13: Continuous manipulation of the Wavy Hair attribute. Each row shows a smooth transition with identity preserved.
Figure 14: Continuous manipulation of the Wearing Lipstick attribute. Each row shows a smooth transition with identity preserved.
Figure 15: Continuous manipulation of the Afro hairstyle attribute. Each row shows increasing Afro intensity from left to right, with identity and background preserved.
Figure 16: Continuous manipulation of the Bald attribute. Each row shows increasing Bald intensity from left to right, with identity and background preserved.
Figure 17: Manipulation for the Blond Hair attribute. Our method alters hair color while maintaining identity and realism.
Figure 18: Manipulation of the Eyeglasses attribute. Each pair shows the original image and its edited version with eyeglasses added.
Figure 19: Attribute manipulation results for Hat. Each pair shows the subject before and after adding a hat.
@@
Figure 20: Attribute manipulation results across various appearance traits. Each pair shows the subject before (left) and after (right) editing, demonstrating the effect of adding attributes such as black hair, scar, necklace, beard, receding hairline, and multicolored hair. The results illustrate the model’s ability to generate realistic and identity-consistent modifications across a diverse set of visual changes.

### v2

@@ B.1 Additional attribute Manipulation
We present additional qualitative results demonstrating our method’s ability to manipulate a wide range of attributes in an identity-preserving manner, as illustrated in Figures 16 to 19. Each figure below shows a series of faces in rows, where the leftmost column is the original input image, and the subsequent columns represent progressive edits along a specific semantic attribute.
B.2 Training SAE with FLUX
To evaluate the generalization of our method, we train the SAE on the FLUX.1-dev [26] model. Text embeddings are extracted from the 23rd layer of the T5 text encoder (out of 24 layers in total). Qualitative results of this experiment are presented in Figure 13.
Figure 13: Smile manipulation on Flux using our method. The generated results maintain the original identity and scene context while achieving realistic smile steering.
B.3 Comparison with Raw Embeddings
We compare our sparse direction, $W_{\text{dec}}(\lambda\times\text{ENC}(x_{\text{A}}))$, against raw embedding addition, $\lambda\times x_{\text{A}}$, in Table 3. Our method consistently outperforms $\text{Emb}_{\text{raw}}$ across all attributes and metrics, yielding an average improvement of 0.212 in QS and 0.196 in IS. This confirms that SAE-derived directions capture purer semantic signals than the raw embeddings from the pretrained text encoder. As noted in [3], original embedding spaces are often dense and semantically entangled, which can cause unintended identity shifts (see Fig.1 in the main paper). By mapping these into a high-dimensional sparse space, our approach achieves semantic disentanglement by decomposing representations into their underlying, interpretable components. This alignment with SpLICE [3] further demonstrates the effectiveness of decomposing dense representations into sparse, clean units.
Table 3: Comparison with raw attribute embedding.
Method Old Smile Makeup Avg
QS IS QS IS QS IS QS $\uparrow$ IS $\uparrow$
[…]
Ours 4.049 0.716 4.265 0.637 4.291 0.742 4.202 0.698
Figure 14: Geometric edit continuity and linearity comparison.
B.4 Geometric metrics for continuity
We adopt dlib landmarks to track facial geometry changes, computing L2 pixel distances within the edit region across steering strengths. As shown in Figure 14, all methods achieve monotonic geometric progression, with our method exhibiting a moderate edit magnitude (neither over-editing like CSlider nor under-editing like AttControl). More importantly, our method achieves the highest linearity ($R^{2}$ of 0.973), compared to CSlider (0.966) and AttControl (0.962), confirming its superior precision in continuous control.
B.5 Visualize sparsity of latent space
We compare raw activations (i.e., CLIP text embeddings) with those in our SAE latent space. As shown in Figure 15, raw embeddings are semantically entangled, whereas our SAE yields a sparse, axis-aligned representation. In this space, distinct dimensions are selectively triggered for specific attributes (Smile vs.Unsmile) samples.
Figure 15: Raw vs. SAE latent space sparsity.
Appendix C Prompt Information
@@
For reproducibility, we provide the text prompts used to generate the images in the appendix figures.
C.1 Prompts for Main Figures
@@
The text prompts used to generate the images for the main figures are as follows.
@@
Figure 1: (1) A photo of a gentle Thai man with layered hair, in a warm-toned studio.
@@
Figure 1: (2) A photo of an honest Dutch man in a white button-up shirt, with a blurred background.
@@
Figure 1: (3) A high-quality portrait of a man in a tuxedo, with clear, detailed facial features.
@@
Figure 4: Prompt: A close-up of a woman wearing a red scarf, looking thoughtful.
Figure 5: A photo of an honest Latin American talented woman.
@@
Figure 6: (1) A woman with medium-length blonde hair, pale skin, and soft blue eyes.
Figure 6: (2) A photo of a peaceful Middle Eastern woman.
Figure 7: A photo of a woman wearing a red scarf.
@@
Figure 9: (1) A photo of a forest.
Figure 9: (2) A path through a dense forest.
Figure 10: (1) A portrait of a woman and a man, standing under cherry blossoms.
Figure 10: (2) A portrait of a woman and a man holding cameras in a motorcycle garage with subtle tones.
Figure 10: (3) A portrait of a woman and a man, wearing scarves in a winter city, in gentle morning sunshine.
Figure 10: (4) A portrait of a woman and a man wearing photography gear at a music studio in a cool color palette.
C.2 Prompts for Appendix Figures
The text prompts used to generate the images in the appendix are listed below.
@@ C.2 Prompts for Appendix Figures
Smile: (1) A photo of a charming Singaporean man with clean-cut hair, with studio backdrop.
(2) A photo of a composed Moroccan man wearing a woolen scarf, with a blurred background.
(3) A photo of an honest Dutch man in a white button-up shirt, with a blurred background.
Old: (1) A photo of a refined Italian man wearing a checkered blazer, with soft window light.
(2) A photo of a serene Danish man with round glasses, near a sunlit window.
(3) A photo of a poised Greek person in a linen shirt, in a studio with soft shadows.
Wearing-Lipstick: (1) A photo of an honest Latin American talented woman.
(2) A woman with short hair, looking at the camera.
Afro
: A portrait of a woman standing by the beach at sunset, with soft and natural lighting.
Wavy-Hair
: A photo of a composed Swedish man with short blond hair, with a blurred background.
@@
Eyeglasses: (1) A photo of an elegant Chinese man wearing a red scarf, with a blurred background.
Eyeglasses
: (2) A photo of a cheerful Filipino man wearing a red scarf, with a soft gray backdrop.
Hat
: (1) A photo of a beautiful woman.
Hat
: (2) A photo of a wise European man wearing a casual sweater, in a sunny environment.
Bald
: (1) A photo of a gentle Vietnamese man with short cropped hair, with a soft gray backdrop.
Bald: (2) A photo of a serene Danish man with round glasses, near a sunlit window.
Blond-Hair
: (1) A photo of a sophisticated Russian woman with a pearl necklace, in a minimalist room.
Blond-Hair
: (2) A woman with short hair, looking at the camera.
Figure 18 photography style prompts:
(1) A photo of sun rays through tall trees.
(2) A peaceful scene of a lake at sunset.
(3) & (6) A group of tall trees in a park.
(4) & (5) A photo of a lone tree in a vast field.
(7) A sunrise over a calm ocean.
(8) A photo of a snowy forest path.
@@
(1) A portrait of a woman and a man framing a shot together in a snowy village in cinematic lighting.
(2) A close-up of a woman and a man, front of a cafe, with crisp autumn light.
(3) A portrait of a woman and a man, wearing scarves in a winter city, in gentle morning sunshine.
@@
(4) A portrait of a woman and a man standing by the beach at sunset, with soft and natural lighting.
@@
Figure 16: Continuous manipulation of the Smile, Old, Wearing-Lipstick, Afro and WavyHair attribute. Each row shows a subject transitioning from neutral to the attribute steered, with identity and lighting well preserved.
@@
Figure 17: Attribute manipulation results across various appearance traits. Each pair shows the subject before (left) and after (right) editing, demonstrating the effect of adding attributes, such as Eyeglasses, Hat, Bald, and BlondHair. The results illustrate the model’s ability to generate realistic and identity-consistent modifications across a diverse set of visual changes.
Figure 18: Continuous manipulation of photography style, such as the Black-and-White, Golden-Hour, Pastel-Tone and Neon-Lighting attribute. Each row demonstrates the style transition on the original images while maintaining their core structural consistency.
Figure 19: Continuous manipulation of attribute in a multi-subject scene. Each row demonstrates the continuous application of an attribute (e.g., Smile, Old) to a specific subject (e.g., man or woman) while keeping the other subject unchanged.